OpenStudy (anonymous):

find then inverse of h if h(x)=-8x/x-4. state the domain and range

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

waht

OpenStudy (anonymous):

\[\frac{-8x}{x -4} =h(x)\]

OpenStudy (anonymous):

Lets take h(x) = y

OpenStudy (anonymous):

-8x = y x - 4y

OpenStudy (anonymous):

8x + yx = 4y \[x = \frac{4y}{8 + y}\]

OpenStudy (anonymous):

This is your Inverse \[h(x) = \frac{4x}{8 + x}\]

OpenStudy (anonymous):

Now domain should real numbers except x = -8

OpenStudy (anonymous):

(-inf,-8)U(-8,inf)?

OpenStudy (anonymous):

As domain and (-inf,4)U(4,inf) as range?

OpenStudy (anonymous):

Yeah You can say that too

OpenStudy (anonymous):

haha I just want to check with you on everything :D

OpenStudy (anonymous):

sharks is right for domain and range too